State the domain of the given function.
f(x) = (2)/(2x-5)
A) (5/2,∞)
B) (-∞, ∞)
C) all real numbers except (5/2)
D) all real numbers except (2/5)

Answer by RAY100(1637)

f(x) = x/ (2x-5)
.
function is valid except when (2x-5) =0,,or 2x=5, or x=5/2
.
Domain is all Real Numbers Except x=5/2,, which is answer C